OpenStudy (anonymous):

(-4,2)(-8,-3) Put this in slope intercept form and standard form

OpenStudy (anonymous):

m=(y1-y)/(x1-x) m=(2-(-3))/(-4-(-8))

OpenStudy (anonymous):

m=5/4 y-2=5/4 (x-(-4) y=2+5/4 x +4*5/4 y=5/4 x +7

OpenStudy (anonymous):

inik give you slope intercept form , you need standard form?